OpenStudy (anonymous):

find the smallest number by which 9408 must be divided so that it becomes a perfect square so obtained

OpenStudy (anonymous):

its dam urgent i need to send the solution to my couzin

OpenStudy (anonymous):

@jackellyn

OpenStudy (jackellyn):

9408 = 2 × 2 × 2 × 2 × 2 × 2 × 3 × 7 × 7 If we divide 9408 by 3, then 9408 ÷ 3 = 3136 = 2 × 2 × 2 × 2 × 2 × 2 × 7 × 7, which is a perfect square. Therefore, the required smallest number is 3. Hence, 3136 = 2 × 2 × 2 × 7 = 56 I found this online, @ http://bit.ly/LCgaLW

OpenStudy (anonymous):

thank you so much...........

OpenStudy (anonymous):

@jackellyn so basically the answer to this question is 3

OpenStudy (anonymous):

?

OpenStudy (jackellyn):

The answer should be 3 @nitz unless they also ask for the perfect square also.